OpenStudy (anonymous):

Find the y-intercept of x-2y=4 and some of my answers are (-2,0) , (0,-2), (0,4) ,(4,0)

OpenStudy (anonymous):

the y intercept would be (0,-2)

OpenStudy (the_fizicx99):

To find the y intercept set x for 0 To find for x intercepts set y for 0 So, x-2y=4 to find y set x for 0 0 -2y = 4 -2y is the y intercept or (0,-2)

OpenStudy (anonymous):

Let's try (-2, 0). Plug in the numbers in that equation. x - 2y =4 Which would be.. -2 - 2(0) = 4 = -2 - 0 = 4 = -2 How about (0, -2). Again plug in the numbers into the equation. x - 2y = 4 That is.. 0 - -2(-2) = 4 = 0 - 4 = 4 = -4 = 4... Um.... I tried my best on that. I know it ISSS (0, -2). Maybe.. Switch the numbers? I don't know. I am not very good at explaining things.